# IF-Critic: Towards a Fine-Grained LLM Critic for Instruction-Following Evaluation

### v1

@@ Abstract
Instruction following is a fundamental ability of Large Language Models (LLMs), requiring their generated outputs to follow multiple constraints imposed in input instructions. Numerous studies have attempted to enhance this ability through preference optimization or reinforcement learning based on reward signals from LLM-as-a-Judge. However, existing evaluation models for instruction following still possess many deficiencies, such as substantial costs and unreliable assessments. To this end, we propose IF-Critic, an LLM critic that can provide efficient and reliable assessments of constraint following in the instructions. We first develop a checklist generator to decompose instructions and generate constraint checklists. With the assistance of the checklists, we collect high-quality critique training data through a multi-stage critique filtering mechanism and employ a constraint-level preference optimization method to train IF-Critic. Extensive experiments demonstrate that the evaluation performance of IF-Critic can beat strong LLM-as-a-Judge baselines, including Deepseek-R1 and o4-mini. With the scalable reward signals provided by IF-Critic, LLMs can achieve substantial performance gains in instruction-following optimization under lower computational overhead compared to strong LLM critic baselines.
@@
Work done when this author interned at Zhipu AI.Equal contribution.Corresponding author.
Codehttps://github.com/thu-coai/IF-CRITIC
Introduction
@@
Large language models (LLMs) have demonstrated remarkable capabilities in various NLP tasks (Zhao et al. 2023). Among these, instruction following stands out as one of the most crucial requirements for LLM applications (Ouyang et al. 2022). In the real-world use of LLMs, nearly all tasks are formulated as instruction following, where human instructions specify task requirements and impose corresponding constraints on the model output (Jiang et al. 2024). Accurately following instructions serves as a key factor in ensuring the helpfulness and reliability of LLMs (Huang et al. 2024).
@@
However, LLMs still have not excelled in instruction following, especially when dealing with complex instructions that involve numerous constraints (Jiang et al. 2024; Wen et al. 2024). Numerous research attempts have been made to improve the instruction-following ability of LLMs. They often leverage LLMs to synthesize complex instructions and perform preference optimization or reinforcement learning based on the reward signals from LLM-as-a-Judge (Sun et al. 2024; Zhang et al. 2024c; He et al. 2024a; Ren et al. 2025; Liu et al. 2025b; Peng et al. 2025).
@@
Figure 1: A usage example of IF-Critic: Given an instruction and a response, a checklist generator first decomposes the instruction to generate a constraint checklist. Then, IF-Critic can provide fine-grained evaluations for the response with respect to its following of all included constraints.
@@
Although these works have achieved promising results, the usage of LLM-as-a-Judge for instruction-following still faces two major yet underexplored challenges: (1) Substantial Costs: Existing methods typically rely on proprietary LLMs (e.g., GPT-4o) or large reasoning models (e.g., QwQ-32B) to provide evaluation results as reward signals, which is costly and difficult to scale up in training data construction. (2) Unreliable Assessments: Current LLMs still struggle to accurately assess instruction-following (Zhang et al. 2024b), exhibiting low recall in error detection (Kamoi et al. 2024) and limited counting ability (Zhang, Cao, and You 2024; Fu et al. 2024). These limitations introduce noise to the evaluation and restrict the gains in model optimization. To tackle the above challenges, some works attempt to introduce handcrafted constraints that can be automatically verified by code execution in data construction, aiming to enhance the efficiency and reliability of evaluation (Dong et al. 2025; Kim et al. 2025; Liu et al. 2025b; Peng et al. 2025). However, these constraints are often homogeneous and atomized, failing to cover the diversity and complexity of human instructions, such as constraints composition (Wen et al. 2024).
@@
In this work, we propose IF-Critic, an LLM critic with 14B parameters tailored for instruction-following evaluation. To enhance inference efficiency and obtain a more holistic and granular perception of the instructions, IF-Critic adopts a novel checklist-guided critique generation paradigm. As shown in Figure 1, we first utilize a checklist generator to decompose the instructions and generate constraint checklists. Then, IF-Critic can evaluate the following of all included constraints in one inference pass, eliminating the need to generate separate judgments for different constraints multiple times. To train IF-Critic, we elaborately prompt Deepseek-R1 (Guo et al. 2025a) to generate expert critiques guided by the checklists, then employ a multi-stage critique filtering mechanism to select the highest-quality critique for each constraint, which mitigates noise and significantly enhances the reliability of critiques for various constraints, including those involving counting. After fine-tuning on this data, we leverage the property that the critiques contain multi-constraint evaluation results in the checklist, further improving IF-Critic via a novel constraint-level preference optimization method. This method performs fine-grained, constraint-level comparisons between positive and negative critiques using Direct Preference Optimization (DPO) (Rafailov et al. 2023), reinforcing the perception of crucial preference information and leading to better alignment between IF-Critic and the expert critiques. Empirically, IF-Critic consistently outperforms various LLM-as-a-Judge baselines on four instruction-following meta-evaluation benchmarks, including the latest state-of-the-art LLMs Deepseek-R1 and o4-mini (Jaech et al. 2024).
@@
Leveraging the fine-grained evaluation results from IF-Critic as reward signals, we adapt DPO and Group Relative Policy Optimization (GRPO) (Shao et al. 2024) to improve the instruction-following ability of LLMs. Experiments demonstrate that IF-Critic brings substantially greater improvements over strong baselines with significantly less computational overhead, highlighting its effectiveness and practicability. Our contributions can be summarized as follows:
@@
  • We introduce IF-Critic, the first instruction-following critic with 14B parameters, developed through our framework. Experiments show that IF-Critic achieves superior evaluation performance compared to strong LLM-as-a-Judge baselines, including Deepseek-R1 and o4-mini.
  • We validate the potential of IF-Critic to provide scalable reward signals and improve the instruction-following abilities of LLMs using DPO and GRPO. IF-Critic can bring substantially greater improvements under lower computational overhead compared to strong baselines.
Related Work
@@
Utilizing LLMs as text evaluators has gradually become prevalent (Wang et al. 2023a; Chen et al. 2023; Ke et al. 2023) due to their advantages in flexibility, interpretability, and generalization. Although proprietary LLMs have demonstrated excellent evaluation capabilities (Zheng et al. 2023; Wang et al. 2023a), their practical application is limited by issues such as high cost and potential data leakage (Ke et al. 2024). To reduce the reliance on proprietary LLMs, a series of works attempt to train smaller LLMs specialized in evaluation using the data generated from humans or proprietary LLMs (Kim et al. 2024a; Li et al. 2024; Ke et al. 2024; Hu et al. 2024; Kim et al. 2024b). Distinct from these works that focus on evaluating overall text quality, we develop an evaluation model that provides constraint-level, fine-grained assessments of instruction-following and leverages them as reward signals to enhance the instruction-following ability of LLMs.
Instruction Following Optimization.
Instruction following remains a critical factor in determining the practicality of LLMs (Liu et al. 2023; Lou, Zhang, and Yin 2024), driving extensive studies to enhance this ability. Earlier works primarily focus on following simple instructions that mostly consist of a single constraint (Chiang et al. 2023; Wang et al. 2023c). With the expanding application of LLMs, recent studies place greater emphasis on complex instruction-following that involves numerous constraints. They typically leverage LLMs to synthesize complex instructions and perform preference optimization or reinforcement learning based on the reward signals from LLM-as-a-Judge (Sun et al. 2024; Zhang et al. 2024c; Cheng et al. 2025; Ren et al. 2025; Liu et al. 2025b; Peng et al. 2025). Specifically, SPaR (Cheng et al. 2025) employs self-play tree search refinement through a fine-tuned refiner to construct DPO training data. RECAST (Liu et al. 2025b) categorizes constraint synthesis into soft and hard constraints, leveraging GPT-4o and code execution, respectively, to obtain evaluation results for GRPO training. However, the effectiveness of existing methods is restricted by the high costs and limited instruction-following evaluation ability of LLM-as-a-Judge, and code-verifiable constraints fail to cover the diversity and complexity of human instructions.
IF-Critic Framework
@@
Task Definition and Method Overview
Distinct from previous LLM critics that typically generate evaluation results for a single criterion or question (Wang et al. 2024a; Vu et al. 2024), IF-Critic leverages a constraint checklist to guide critique generation and evaluates the following of all constraints included in the checklist within one inference pass. Specifically, given an instruction $x$, corresponding response $y$, and a constraint checklist $\{c_{k}\}^{n}_{k=1}$ which contains all constraints within $x$, IF-Critic can provide a critique $C=\bigcup_{k=1}^{n}(e_{k},j_{k})$ for $y$, where each segment $(e_{k},j_{k})$ consists of an explanation $e_{k}$ and a binary judgment $j_{k}$ (0 or 1) for the respective constraint following. This paradigm not only enhances inference efficiency but also equips IF-Critic with a more holistic and granular perception of the instructions, thus improving the reliability of instruction-following assessments.
@@
Instruction and Response Collection
To obtain diverse instructions for critic training, we collect instructions from real application scenarios and utilize LLMs to categorize them according to the task taxonomy of CritiqueLLM (Ke et al. 2024), which encompasses 10 categories and covers diverse NLP applications in real-world scenarios. To improve data quality, we prompt LLMs to score instruction quality and develop a classifier to assess their constraint complexity, which is trained with a small amount of human annotation data. Finally, we balance the amount of instruction across task categories and acquire 55K high-quality instructions with relatively high complexity.
Then, we collect LLM responses from 15 representative models, which possess varying abilities in instruction-following, including GPT-4o (OpenAI 2023), GPT-4o-mini, seven versions of GLM-4 (GLM et al. 2024), three versions of Qwen2.5 (Yang et al. 2024), Deepseek-R1 (Guo et al. 2025a), Doubao1.5 (Doubao Team 2025), and Claude3.7-sonnet (Anthropic 2023). For each instruction, we randomly select two models to generate responses, resulting in a total of 110k evaluation inputs, each comprising an instruction and its corresponding LLM response.
Critique Training Data Construction
To assist IF-Critic in providing fine-grained assessments of instruction-following, we first develop a checklist generator to decompose the constraints within instruction $x$ and generate a constraint checklist $\{c_{k}\}^{n}_{k=1}$. We leverage a powerful LLM, Deepseek-R1111https://huggingface.co/deepseek-ai/DeepSeek-R1, for automatic checklist annotation in our collected instructions and utilize its generated checklists to fine-tune the checklist generator. Manual inspection of 1,000 test samples shows the accuracy of our checklist generator exceeds 99%, indicating its reliability.
With the assistance of the checklists, we proceed to employ Deepseek-R1 for expert critique annotation via a carefully designed prompt, which elicits a concise and specific explanation before the judgment for each constraint. Detailed prompts are in the Appendix. For each response, we collect $N$ critiques and adapt a multi-stage critique filtering mechanism to select the highest-quality critique for each constraint in the checklist, which is then used for training IF-Critic.
The mechanism comprises four stages. Specifically, we first employ cross-model and rule-augmented verification to filter out noise in critiques caused by two major limitations of LLM-as-a-Judge: potential evaluation bias and limited counting ability, respectively. Subsequently, we select the highest-quality judgements and explanations from the resulting noise-free critiques as the final critiques based on consistency, significantly enhancing their reliability.
@@
Previous works point out that LLM-as-a-Judge may suffer from potential biases (Zheng et al. 2023; Ye et al. 2025a), undermining the reliability of its generated critiques. To mitigate this, we utilize two additional LLMs, GLM-4-Plus and Qwen2.5-72B-Instruct, to independently verify two aspects of the expert critiques for each constraint: (1) The correctness of the explanation, and (2) The consistency between the explanation and judgment. Detailed prompts can be found in the Appendix. Any critiques that fail verification by any model in any aspect are strictly filtered, accounting for 11.3% of our data.
@@
Considering that LLMs often struggle with accurate counting (Zhang, Cao, and You 2024; Fu et al. 2024), we introduce rule-augmented verification to enhance the reliability of critiques involving length constraints (e.g., each title must be 8 characters long). Methodologically, we first prompt Qwen2.5-72B-Instruct to identify length constraints and extract the segments from the responses that are relevant to these constraints. Subsequently, we employ Python scripts to automatically identify length information within these segments, and require Deepseek-R1 to revise its critiques with the assistance of this information. The prompts for these steps are in the Appendix. This approach integrates the flexibility of LLM-based evaluation with the precision of rule-based evaluation, thereby effectively handling the diverse length constraints in human instructions.
@@
The correctness of judgment is vitally important in instruction-following evaluation. Given that critique generation is analogous to a Chain-of-Thought (CoT) (Wei et al. 2022) reasoning process, we incorporate self-consistency (Wang et al. 2023b), a widely used decoding strategy in CoT reasoning, to select final judgments and improve their correctness. Specifically, for each constraint $c_{k}$, we select the final judgment $j_{k}^{*}$ via majority voting among multiple expert critiques. Furthermore, previous works demonstrate that confidence, which is calculated as the proportion of the majority result to the total number of samplings, can serve as an important indicator of answer reliability (Xiong et al. 2024; Hu et al. 2024; Prasad et al. 2024), with higher confidence generally implying greater reliability. Inspired by this, we discard the final critiques of constraints with confidence lower than 0.75 in our experiment, which empirically balances data retention and reliability.
@@
After obtaining the correct final judgments, the last step is to select the highest-quality explanations that support these judgments. To achieve this, we devise a selection strategy inspired by Minimum Bayes Risk Decoding (Bickel and Doksum 2015). Specifically, for each constraint $c_{k}$, we first construct a hypothesis set $\mathcal{H}_{k}$ containing all explanations whose corresponding judgments align with the final judgments $j_{k}^{*}$. Then, we utilize a text similarity measure $u(\cdot)$ to calculate the consistency between each pair of explanations in $\mathcal{H}_{k}$. The explanation with the highest average consistency among others is selected as the final explanation $e_{k}^{*}$. The formal definition is as follows:
@@
IF-Critic Training
After constructing the final critique training dataset $D=\{C_{i}=\bigcup_{k=1}^{n_{i}}(e_{ik}^{*},j_{ik}^{*})\}_{i=1}^{|D|}$, we partition $D$ into two subsets $D_{\text{sft}}\cup D_{\text{ref}}$ and employ a two-stage paradigm to train IF-Critic, comprising supervised fine-tuning (SFT) and constraint-level preference optimization. In the first stage, we utilize a critique generation prompt template $p$ to concatenate the instruction $x$, response $y$, and checklist $\{c_{k}\}^{n}_{k=1}$, applying a standard SFT loss on vanilla LLM $\theta$:

### v2

@@ Abstract
Instruction-following is a fundamental ability of Large Language Models (LLMs), requiring their generated outputs to follow multiple constraints imposed in input instructions. Numerous studies have attempted to enhance this ability through preference optimization or reinforcement learning based on reward signals from LLM-as-a-Judge. However, existing evaluation models for instruction-following still possess many deficiencies, such as substantial costs and unreliable assessments. To this end, we propose IF-Critic, an LLM critic for fine-grained, efficient, and reliable instruction-following evaluation. We first develop a checklist generator to decompose instructions and generate constraint checklists. With the assistance of the checklists, we collect high-quality critique training data through a multi-stage critique filtering mechanism and employ a constraint-level preference optimization method to train IF-Critic. Extensive experiments show that the evaluation performance of IF-Critic can beat strong LLM-as-a-Judge baselines, including o4-mini and Gemini-3-Pro. With the reward signals provided by IF-Critic, LLMs can achieve substantial performance gains in instruction-following optimization under lower computational overhead compared to strong LLM critic baselines. Our code and model are available at https://github.com/thu-coai/IF-CRITIC.
IF-Critic: Towards a Fine-Grained LLM Critic for Instruction-Following Evaluation
Bosi Wen1,†,∗   Yilin Niu2,∗  Cunxiang Wang2  Pei Ke3  Xiaoying Ling2 Ying Zhang2  Aohan Zeng4  Hongning Wang1   Minlie Huang1,‡ 1The Conversational Artificial Intelligence (CoAI) Group, Tsinghua University 2Zhipu AI  3University of Electronic Science and Technology of China 4The Knowledge Engineering Group (KEG), Tsinghua University wbs23@mails.tsinghua.edu.cn, aihuang@tsinghua.edu.cn
@@
Work done when this author interned at Zhipu AI.Equal contributionCorresponding author
1 Introduction
@@
Large language models (LLMs) have demonstrated remarkable capabilities in various NLP tasks Zhao et al. (2023). Among these, instruction-following stands out as one of the most crucial requirements for LLM applications Ouyang et al. (2022). In the real-world use of LLMs, nearly all tasks are formulated as instruction-following, where human instructions specify task requirements and impose corresponding constraints on the model output Jiang et al. (2024). Accurately following instructions serves as a key factor in ensuring the helpfulness and reliability of LLMs Huang et al. (2024).
@@
However, LLMs still have not excelled in instruction-following, especially when dealing with complex instructions that involve numerous constraints Jiang et al. (2024). Numerous research attempts have been made to improve the instruction-following ability of LLMs. They often leverage LLMs to synthesize complex instructions and perform preference optimization or reinforcement learning based on the reward signals from LLM-as-a-Judge Zhang et al. (2025); He et al. (2024a); Ren et al. (2025); Liu et al. (2025b); Peng et al. (2025).
@@
Figure 1: A usage example of IF-Critic: Given an instruction and a response, a checklist generator first decomposes the instruction to generate a constraint checklist. Then, IF-Critic can provide fine-grained evaluations for the response with respect to its following of all included constraints in one inference pass.
@@
Despite these efforts, the usage of LLM-as-a-Judge for instruction-following still faces two major yet underexplored challenges: (1) Substantial Costs: Existing methods typically rely on proprietary LLMs (e.g., GPT-4o) or large reasoning models (e.g., QwQ-32B) to provide evaluation results as reward signals, which is costly and difficult to scale up in model training. (2) Unreliable Assessments: Current LLMs still struggle to accurately assess instruction-following Zhang et al. (2024c), exhibiting low recall in error detection Kamoi et al. (2024) and limited counting ability Zhang et al. (2024b); Fu et al. (2024). These limitations introduce noise to the evaluation and restrict the gains in model optimization. To tackle these challenges, some works attempt to introduce handcrafted code-verifiable constraints in data construction, aiming to enhance evaluation efficiency and reliability Dong et al. (2025); Kim et al. (2025); Li et al. (2025); Peng et al. (2025). However, these constraints are often homogeneous and atomized, failing to cover the diversity and complexity of human instructions, such as constraint composition Wen et al. (2024).
@@
In this work, we propose IF-Critic, an LLM critic tailored for instruction-following evaluation. To enhance inference efficiency and obtain a more holistic and granular perception of the instructions, IF-Critic adopts a novel checklist-guided critique generation paradigm. As shown in Figure 1, we first utilize a checklist generator to decompose the instructions and generate constraint checklists. Then, IF-Critic can evaluate the following of all included constraints in one inference pass, eliminating the need to generate separate judgments for different constraints multiple times. To train IF-Critic, we elaborately prompt Deepseek-R1 Guo et al. (2025a) to generate expert critiques guided by the checklists, then employ a multi-stage critique filtering mechanism to select the highest-quality critique for each constraint, which significantly mitigates noise and enhances the reliability of critiques, including those involving counting. After fine-tuning on this data, we leverage the property that the critiques contain multi-constraint evaluation results, further improving IF-Critic via a novel constraint-level preference optimization method. This method performs fine-grained, constraint-level comparisons between positive and negative critiques using Direct Preference Optimization (DPO) Rafailov et al. (2023), reinforcing the perception of crucial preference information and leading to better alignment between IF-Critic and the expert critiques. Empirically, IF-Critic consistently outperforms various LLM-as-a-Judge baselines on four instruction-following meta-evaluation benchmarks, including the latest state-of-the-art LLMs o4-mini Jaech et al. (2024) and Gemini-3-Pro Team et al. (2023).
@@
Leveraging the fine-grained evaluation results from IF-Critic as reward signals, we adapt DPO and Group Relative Policy Optimization (GRPO) Shao et al. (2024) to improve the instruction-following ability of LLMs. Experiments demonstrate that IF-Critic brings substantially greater improvements over strong baselines with significantly less computational overhead, highlighting its effectiveness and practicability. Our contributions can be summarized as follows:
@@
  • We introduce IF-Critic, the first instruction-following critic developed through our framework. Experiments show that IF-Critic achieves superior evaluation performance compared to strong LLM-as-a-Judge baselines, including o4-mini and Gemini-3-Pro.
  • We validate that IF-Critic can provide scalable reward signals to improve the instruction-following abilities of LLMs using DPO and GRPO. IF-Critic can bring substantially greater improvements under lower computational overhead compared to strong baselines.
2 Related Work
@@
Utilizing LLMs as text evaluators has gradually become prevalent Wang et al. (2023a); Chen et al. (2023); Ke et al. (2023) due to their advantages in flexibility, interpretability, and generalization. Although proprietary LLMs have demonstrated excellent evaluation capabilities Zheng et al. (2023); Wang et al. (2023a), their practical application is limited by issues such as high cost and potential data leakage Ke et al. (2024). To reduce the reliance on proprietary LLMs, a series of works attempt to train smaller LLMs specialized in evaluation using the data generated from humans or proprietary LLMs Kim et al. (2024a); Li et al. (2024); Ke et al. (2024); Hu et al. (2024); Kim et al. (2024b). Distinct from these works that focus on evaluating overall text quality, we develop an evaluation model that provides constraint-level, fine-grained assessments of instruction-following and leverages them as reward signals to enhance the instruction-following ability of LLMs.
instruction-following Optimization.
As LLMs have been increasingly applied to real-world complex tasks, following complex instructions that involve numerous constraints emerges as a key determinant of their practical utility Liu et al. (2023); Lou et al. (2024), driving extensive studies to enhance this ability of LLMs. They typically leverage LLMs to synthesize complex instructions and perform preference optimization or reinforcement learning based on the reward signals from LLM-as-a-Judge Sun et al. (2024); Zhang et al. (2025); Cheng et al. (2025); Ren et al. (2025); Liu et al. (2025b); Peng et al. (2025). Specifically, SPaR Cheng et al. (2025) employs self-play tree search refinement through a fine-tuned refiner to construct DPO training data. RECAST Liu et al. (2025b) categorizes soft and hard constraints, leveraging GPT-4o and code execution, respectively, to obtain evaluation results for GRPO training. However, the effectiveness of existing methods is restricted by the high costs and limited instruction-following evaluation ability of LLM-as-a-Judge, and code-verifiable constraints fail to cover the diversity and complexity of human instructions.
3 Methodology
@@
3.1 Task Definition and Method Overview
Distinct from previous LLM critics that typically generate evaluation results for a single criterion or question Vu et al. (2024); Wang et al. (2025), IF-Critic leverages a constraint checklist to guide critique generation and evaluates the following of all included constraints in one inference pass. Specifically, given an instruction $x$, corresponding response $y$, and a constraint checklist $\{c_{k}\}^{n}_{k=1}$ which contains all constraints within $x$, IF-Critic can provide a critique $C=\bigcup_{k=1}^{n}(e_{k},j_{k})$ for $y$, where each segment $(e_{k},j_{k})$ consists of an explanation $e_{k}$ and a binary judgment $j_{k}$ (0 or 1) for the following of the respective constraint. This paradigm not only enhances inference efficiency but also equips IF-Critic with a more holistic and granular perception of the instructions, thus improving the reliability of instruction-following assessments.
@@ 3.1 Task Definition and Method Overview
3.2 Instruction and Response Collection
To obtain diverse instructions for critic training, we collect instructions from real-world application scenarios and utilize LLMs to categorize them according to the task taxonomy of CritiqueLLM Ke et al. (2024), which encompasses 10 categories and covers diverse NLP applications in real-world scenarios. To improve data quality, we prompt LLMs to score instruction quality and develop a classifier to assess their constraint complexity, which is trained with a small amount of human annotation data. Finally, we balance the amount of instruction across task categories and acquire 55K high-quality instructions with relatively high complexity.
Then, we collect LLM responses from 15 representative models, which possess varying abilities in instruction-following. The model list is provided in Appendix A. For each instruction, we randomly select two models to generate responses, resulting in a total of 110k evaluation inputs, each comprising an instruction and its corresponding LLM response.
3.3 Critique Training Data Construction
To assist IF-Critic in fine-grained evaluation, we first develop a checklist generator to decompose the constraints within instruction $x$ and generate a constraint checklist $\{c_{k}\}^{n}_{k=1}$. We leverage a powerful LLM, Deepseek-R1  [^1] https://huggingface.co/deepseek-ai/DeepSeek-R1, for automatic checklist annotation in our collected instructions. Then, its generated checklists are utilized to fine-tune the generator, which supports efficient checklist generation after deployment. Manual inspection of 1,000 test samples for the checklist generator shows that 99.29% of generated constraints and 97.50% of entire checklists are correct, indicating its reliability.
With the assistance of the checklists, we proceed to employ Deepseek-R1 for expert critique annotation via a carefully designed prompt, which elicits a concise and specific explanation before the judgment for each constraint. Detailed prompts are in the Appendix B. For each response, we collect $N$ expert critiques and adapt a multi-stage critique filtering mechanism to select the highest-quality critique for each constraint in the checklist, which is then used for training IF-Critic.
The mechanism comprises four stages. We first employ cross-model and rule-augmented verification to filter out noise in critiques caused by two major limitations of LLM-as-a-Judge: potential evaluation bias and limited counting ability, respectively. Subsequently, we select the highest-quality judgements and explanations from the resulting noise-free critiques as the final critiques based on consistency, further enhancing their reliability.
@@ 3.3 Critique Training Data Construction
Previous works point out that LLM-as-a-Judge may suffer from potential biases Zheng et al. (2023); Ye et al. (2025a), undermining the reliability of its generated critiques. To mitigate this, we utilize two additional LLMs, GLM-4-Plus and Qwen2.5-72B-Instruct, to independently verify two aspects of the expert critiques for each constraint: (1) The correctness of the explanation, and (2) The consistency between the explanation and judgment. Detailed prompts can be found in Appendix B. Any critiques that fail verification by any model in any aspect are strictly filtered, accounting for 11.3% of our data.
@@
Considering that LLMs often struggle with accurate counting Zhang et al. (2024b); Fu et al. (2024), we introduce rule-augmented verification to enhance the reliability of critiques involving length constraints (e.g., each title must be 8 characters long). Methodologically, we first prompt Qwen2.5-72B-Instruct to identify length constraints and extract relevant segments from the responses. Subsequently, we employ Python scripts to identify length information within these segments automatically. This information then guides Deepseek-R1 to revise its critiques. The prompts for these steps are in the Appendix B. This approach integrates the flexibility of LLM-based evaluation with the precision of rule-based evaluation, thereby effectively handling the diverse length constraints in human instructions.
@@
Given that critique generation is analogous to a Chain-of-Thought (CoT) Wei et al. (2022) reasoning process, we incorporate self-consistency Wang et al. (2023b), a widely used decoding strategy in CoT reasoning, to select final judgments and improve their correctness. Specifically, for each constraint $c_{k}$, we select the final judgment $j_{k}^{*}$ via majority voting among multiple expert critiques. Furthermore, previous works show that confidence, which is calculated as the proportion of the majority result to the total samples, is an important indicator of answer reliability Xiong et al. (2024); Hu et al. (2024); Prasad et al. (2024), with higher confidence generally implying greater reliability. Inspired by this, we discard the final critiques of constraints with confidence lower than 0.75 in our experiment, which empirically balances data retention and reliability.
@@
After obtaining the correct final judgments, the last step is to select the highest-quality explanations that support these judgments. To achieve this, we devise a selection strategy inspired by Minimum Bayes Risk Decoding Bickel and Doksum (2015): For each constraint $c_{k}$, we first construct a hypothesis set $\mathcal{H}_{k}$ containing all explanations whose corresponding judgments align with the final judgments $j_{k}^{*}$. Then, we utilize a text similarity measure $u(\cdot)$ to calculate the consistency between each pair of explanations in $\mathcal{H}_{k}$. The explanation with the highest average consistency among others is selected as the final explanation $e_{k}^{*}$. The formal definition is as follows:
@@
Through these rigorous filtering processes, we manually inspect the quality of 70 final critiques with 353 constraints and find that 96.03% judgments and 92.35% explanations are completely correct, demonstrating their high quality.
3.4 IF-Critic Training
After constructing the final critique training dataset $D=\{C_{i}=\bigcup_{k=1}^{n_{i}}(e_{ik}^{*},j_{ik}^{*})\}_{i=1}^{|D|}$, we partition $D$ into two subsets $D_{\text{sft}}\cup D_{\text{ref}}$ and employ a two-stage paradigm to train IF-Critic, comprising supervised fine-tuning (SFT) and constraint-level preference optimization. In the first stage, we utilize a critique generation prompt $p$ to concatenate the instruction $x$, response $y$, and checklist $\{c_{k}\}^{n}_{k=1}$, applying a standard SFT loss on vanilla LLM $\theta$: